SEARCHING FOR THE LOW-LYING TRIPLET STATE OF CCl_{2}

\underline{ROBERT J. CLARK}^{a}, ANDREW BEZANT, GYÖRGY TARCZAY^{b}, TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, OH 43210; MEI-LUNG LIU, CHIA-LING LEE, BOR-CHEN CHANG, Department of Chemistry, National Central University, Chung-Li 32054, Taiwan.

Dispersed fluorescence spectra have been recorded pumping several Ã − X̃ electronic transition of CCl_{2}. The analysis of the data yield well defined values for the harmonic and anharmonic constants for the symmetric ground state vibrations. Additional lines are observed at red shifts from 5000 cm^{−1} from the origin, which cannot be not assigned to any X̃ vibrational levels. A possible explanation for these transitions would be perturbations of the singlet Ã by the low-lying triplet ã state. This observation is in fair agreement with the theoretical predictions for the singlet-triplet state energy separation, and supports the new, alternative interpretations of the photodetachment spectrum of Lineberger et al.

aPresent address: Department of Chemistry, Ohio Northern University, Ada OH 45810

bPermanent address: Department of General and Inorganic Chemistry, Eötvös University, P.O. Box 32, H-1518 Budapest 112, Hungary